OpenStudy (just_gymnast_me):

Please help! Will fan and medal!!! Solve the proportion x over x minus 4 equals 4 over x minus 4.. What is the value of x? x = 2 x = −2 and x = 1 x = −1 and x = 4 x = 4

OpenStudy (just_gymnast_me):

\[\frac{ x }{ x-4} = \frac{ 4 }{ x-4 }\]

OpenStudy (dan815):

okay well you are dividing by the same number thing so...

OpenStudy (just_gymnast_me):

the numerator x is equal to 4?

OpenStudy (dan815):

yeah

OpenStudy (just_gymnast_me):

then im stuck..is that it?

OpenStudy (dan815):

yep, if u are confused you can plug in the other values and see what happens

OpenStudy (just_gymnast_me):

Ok, thanks (:

OpenStudy (dan815):